OpenStudy (thesecret20111):

Use the pythagorean identity to simplify the following expression: cos^2*50 degrees +cos^2*40 degrees. I think it has something to do with the formula: sin(90-A)=cos A Please help showing steps :)

terenzreignz (terenzreignz):

You're quite right in saying that sin(90 - A) = cos A But it is also true that cos(90 - A) = sin A So go ahead and implement that idea, and replace one of the cosines there with a sine...

OpenStudy (thesecret20111):

I'm not sure how to use the rule, would it be something along the lines of: cos^2(90-50)=sin 40^2?

terenzreignz (terenzreignz):

More like this... \[\huge \cos^2(40^o)=\cos^2(90^o-50^o)=\color{red}?\]

OpenStudy (thesecret20111):

So cos^2*40=cos^2*4=1?

terenzreignz (terenzreignz):

No, see, you use this identity... \[\huge \cos(90^o - A )=\sin(A)\]

terenzreignz (terenzreignz):

By that logic, \[\huge \cos(90^o - 50^o)=\color{red}?\]

OpenStudy (thesecret20111):

cos (90 degrees - 50 degrees)= sin(40)? Sorry if I'm being slow :/

terenzreignz (terenzreignz):

Look more carefully at the parts I emphasize... \[\huge \cos(90^o - \color{blue}A )=\sin(\color{blue}A)\]

OpenStudy (thesecret20111):

so sin(50)?

terenzreignz (terenzreignz):

Yes. And therefore.. \[\huge \cos^2(40^o)=\cos^2(90^o-50^o)=\color{red}?\]

OpenStudy (thesecret20111):

cos^2(40)=sin^2(50)?

terenzreignz (terenzreignz):

Most excellent :) Therefore... cos^2(40) + cos^2(50) = ?

OpenStudy (thesecret20111):

Do I just work out sin^2(50)?

terenzreignz (terenzreignz):

No need. It becomes \[\huge \cos^2(50^o)+\sin^2(50^o)\] right?

OpenStudy (thesecret20111):

Ohh so then I use the rule sin^2A+cos^2A=1?

OpenStudy (agent0smith):

and \[\large \cos^2 \theta + \sin^2 \theta =...?\]

OpenStudy (agent0smith):

Yes :)

terenzreignz (terenzreignz):

Yes, @thesecret20111 precisely :)

OpenStudy (thesecret20111):

Meaning the final answer is 1?

terenzreignz (terenzreignz):

Yup.

OpenStudy (thesecret20111):

Thanks so much! You're awesome! :)

terenzreignz (terenzreignz):

Well, you know, we try :D

OpenStudy (agent0smith):

Which you can also confirm yourself by calculation: http://www.wolframalpha.com/input/?i=cos%5E2%2850%29+%2Bcos%5E2%2840%29

OpenStudy (thesecret20111):

Ok thanks again guys :)